OpenStudy (ayyookyndall):

**WILL MEDAL AND FAN** DUE SOON!!! Given Triangle QRS = Triangle TUV, QS = 3v+2 and TV = 7v-6, find the length of QS and TV

OpenStudy (anonymous):

Is it a scalene/ equilateral triangle?

OpenStudy (ayyookyndall):

There is no picture.

OpenStudy (anonymous):

3v+2=7v-6 -3v -3v 2= 4v-6 +6 +6 8= 4v /4 /4 2=v QS= 3(2)+2= 8 TV= 7(2)-6 = 8